OpenStudy (anonymous):

What is the point-slope form of a line with slope -5 that contains the point (2, -1)? A.y + 1 = 5(x + 2), B.y - 1 = 5(x + 2), C.y + 1 = -5(x - 2), D.y - 1 = -5(x - 2)

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]where \(y_1\) is the y-value, \(m\) is the slope, and \(x_1\) is the x-value so just plug in the given information and your done! :)

OpenStudy (anonymous):

you should end up with \(y+1=-5(x-2)\)

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

as long as you know how to do it now :)

OpenStudy (anonymous):

yes i get it now!!:)

OpenStudy (anonymous):

cool youre welcome!